### v2

# Fast fitting of phylogenetic mixed-effects models

Abstract
Mixed-effects models are among the most commonly used statistical methods for the exploration of multispecies data. In recent years, also Joint Species Distribution Models and Generalized Linear Latent Variale Models have gained in popularity when the goal is to incorporate residual covariation between species that cannot be explained due to measured environmental covariates. Few software implementations of such models exist that can additionally incorporate phylogenetic information, and those that exist tend to utilize Markov chain Monte Carlo methods for estimation, so that model fitting takes a long time. In this article we develop new methods for quickly and flexibly fitting phylogenetic mixed-effects models, potentially incorporating residual covariation between species using latent variables, with the possibility to estimate the strength of phylogenetic structuring in species responses per environmental covariate, and while incorporating correlation between different covariate effects. By combining Variational approximations, a sparse approximation to the phylogenetic precision matrix, and parallel computation, phylogenetic mixed-effects models can be fitted much more quickly than the current state-of-the-art. Two simulation studies demonstrate that the proposed combination of approximations is fast and enjoys high accuracy. We explore sensitivty of the approximation to the ordering of species with a real world dataset of wood-decaying fungi.
Word count: $1.1172\times 10^{4}$
Introduction
The exploration of species co-occurrence patterns is at the core of contemporary community ecology. Recent years have seen a push in methods that better facilitate ecologists to understand which species co-occur, and why they co-occur (Ives and Helmus 2011; Braga et al. 2018). Some species co-occur because they thrive in similar environments, which may be due to similar characteristics in terms of functional traits. Species that share evolutionary history often share such physical characteristics, and consequently hypotheses exist that describe the co-occurrence of closely related species (Pagel 1999).
Data on ecological communities is increasingly analysed with multivariate statistical models. Joint Species Distribution Models (JSDMs, Pollock et al. 2014) are now frequently applied for the analysis of binary data of multiple species. JSDMs are a type of Generalized Linear Mixed-effects Model (GLMM, Brooks et al. 2017; Bates et al. 2015) that incorporates residual covariation to better represent species’ co-occurrence patterns. Generalized Linear Latent Variable Models (GLLVMs, Warton et al. 2015) are a modeling framework that leverages dimension reduction methods to more efficiently implement such multispecies models. GLLVMs have also been presented as a model-based method for ordination (Hui et al. 2015; van der Veen et al. 2023), and have been at the forefront of computational advances for quickly fitting statistical models with random effects (Niku et al. 2019; Korhonen et al. 2021; Kidzinski et al. 2022).
Multivariate statistical models tend to be challenging to fit with species that do not have many observations. Such species may be referred to as “rare” by ecologists, but we here more generally refer to species that have few observations, such as due to insufficient sampling. In order to ameliorate such issues with few data points, and to incorporate the relatedness of species into the analysis of species co-occurrence patterns, JSDMs have been developed that phylogenetically structure species responses to the environment. In such models, species that are more closely related can be predicted to co-occur (Ovaskainen et al. 2017). This has the benefit of “sharing” information across related species, so that species may more accurately be placed in the environment. In fact, with such models it is possible to predict the environmental preferences of extinct species for which no response observations are available at all, by utilizing phylogenetic information.
Three notable software implementations for such phylogenetic mixed-effects models are the Hmsc (Tikhonov et al. 2024), MCMCglmm (Hadfield 2010), and phyr (Ives et al. 2020) R-packages. Unfortunately, these software implementations are either slow because they use Markov chain Monte Carlo for estimation (Hmsc and MCMCglmm) or are limited in their functionality otherwise (e.g., phyr supports few datatypes). Since phylogenetic information and community data of species is becoming more readily available thanks to various openscience initiatives [Hinchliff et al. (2015);], this motivates development of better, and faster, software implementations of community models that incorporate phylogenetic information.
In this article, we present a new statistical approach for fast fitting of phylogenetic mixed-effects models for community ecological data as part of the gllvm R-package (Niku et al. 2023). Phylogenetic information can facilitate better estimation of species responses to the environment, while latent variables can potentially be included to account for residual covariation. The proposed modeling framework can alternatively be used for when response data are functional traits, or data on individuals (with a pedigree instead of a phylogeny as in MCMCglmm), but we choose to focus on community data in this article. The software implementation supports a range of datatypes commonly found in community ecology, such as for counts of individuals (Poisson and negative-binomial, potentially zero-inflated), cover classes (ordinal), percentage cover (beta models potentially with zeros or ones, Korhonen et al. 2024), and biomass (Tweedie models, Korhonen et al. 2021; Niku et al. 2017). Niku et al. (2021) implemented species-specific random effects as part of the fourth corner latent variable model. We extend their work by; 1) phylogenetically structuring the random effects, 2) for which we provide functionality to determine strength of the phylogenetic signal per covariate, and 3) facilitate specification of correlation between the random effects for different covariates using the user-friendly lme4 formula interface (Bates et al. 2015, 2025)
Using a combination of methodological developments we can fit the models much more quickly than existing software implementations for phylogenetic mixed-effects models. Variational Approximations (VA, Ormerod and Wand 2010) helps to retrieve a closed form approximation to the likelihood, which we combine with a matrix normal covariance structure for the variational distribution, as well as a reduced-rank approximation to further speed it up. A sparse approximation to the phylogenetic precision matrix, in the form of a band pattern (Wist and Rue 2006) or with Nearest Neighbour Gaussian processes (NNGPs, Finley et al. 2019) facilitates even quicker computation, i.e., by fitting the models using Gaussian Markov Random fields. Recently, Matsuba et al. (2024) also employed NNGPs to quickly and accurately retrieve the inverse of the phylogenetic covariance matrix, although in a Bayesian context and following for a different model formulation. Finally, we fit the models in parallel via the TMB R-package (Kristensen et al. 2015) to further reduce computation time.
With extensive simulations we show that the proposed combination of four approximations still allows us to accurately retrieve the model parameters, and that it tends to be even faster than fitting the models with the Laplace approximation (e.g., as used in the phyr R-package). The approximation error for the phylogenetic precision matrix is sensitive to row-column permutation (Guinness 2018), an issue that we explore using real data on wood-decaying fungi from Abrego et al. (2022a).
Model formulation
In the following sections we use superscripts to distinguish different quantities, subscripts for indices, bold capitalized characters for matrices, bold lowercase characters for vectors, and ordinary characters to denote scalars. In this section we present and explain the phylogenetic mixed-effects model, and in the next section we discuss estimation of the parameters.
Let Y represent a multivariate dataset with observations $y_{ij}$, where the rows $i=1,\ldots,n$ represent the sites at which an ecological community has been surveyed, and the columns $j=1,\ldots,m$ represent the species found. Let $\textbf{X}=\begin{pmatrix}\textbf{x}_{1},\textbf{x}_{2},\ldots,\textbf{x}_{k}\end{pmatrix}$ denote a set of $k=1,\ldots,p$ covariates, where $\textbf{x}_{i}=\begin{pmatrix}x_{i1},x_{i2},\ldots x_{nk}\end{pmatrix}^{\top}$ with an associated $p\times m$ coefficient matrix B and entries $\beta_{jk}$ for the environment that has been observed at each site, usually including an intercept column. Covariates can be categorical to represent species-specific intercepts under different conditions, or numerical to represent species-specific responses to the environment, i.e., as slope parameters. As in generalised linear models (GLMs) a link function $\text{g}(\cdot)$ connects the model to the conditional mean of the distribution assumed for the responses, so that the model generically is
[…]
This model can additionally be extended with functional traits as in the fourth corner latent variable model (Niku et al. 2021), by hierarchically modeling the species responses
$\textbf{B}=\textbf{1}^{\top}\boldsymbol{\beta}^{x}+\textbf{T}\textbf{B}^{tx}+\textbf{B}^{\epsilon},$
where 1 is a $m$-sized vector of ones, $\boldsymbol{\beta}^{x}=\begin{pmatrix}\beta^{x}_{1},\beta_{2},\ldots,\beta^{x}_{p}\end{pmatrix}^{\top}$ is a $p$-sized vector of community mean responses to the covariates, T is a $m\times t$ matrix of species covariates (most commonly functional traits), $\textbf{B}^{tx}$ is a $t\times p$ matrix of interaction coefficients with columns $\boldsymbol{\beta}^{tx}_{k}$, and $\textbf{B}^{\epsilon}$ is a matrix of $p\times m$ species-specific random effects for the covariates that model species’ deviations from the community mean response with entries $\beta^{\epsilon}_{jk}$. The B with coefficients can be thought of as latent traits, which are modeled with species covariates. Excluding the species covariates thus simplifies the model to only include “residual” information.
Let C denote the phylogenetic correlation matrix of dimension $m\times m$, and $\boldsymbol{\mu}_{k}=\boldsymbol{\beta}^{x}_{k}+\textbf{T}\boldsymbol{\beta}^{tx}_{k}$ the community mean response to the $k^{th}$ covariate. To phylogenetically structure the random effects we could assume a model with a single set of random effects $\textbf{b}^{a}_{k}\sim\mathcal{N}(\boldsymbol{\mu}_{k},\gamma_{k}^{2}\textbf{C})$, so that the variance of the random effect $\gamma_{k}^{2}$ represents the rate of evolution. However, in the absence of a phylogenetic signal this model is quite restrictive, so phylogenetic random effects models are usually formulated with a second set of random effects that are independent for all species, $\textbf{b}^{2}_{k}$ with zero mean and variance $\delta_{k}^{2}$. Since $\textbf{b}^{a}_{k}$ and $\textbf{b}^{b}_{k}$ are independent, it follows that $\boldsymbol{\beta}_{k}=\textbf{b}^{a}_{k}+\textbf{b}^{b}_{k}\sim\mathcal{N}(\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})$ where $\boldsymbol{\Sigma}_{k}=\sigma^{2}_{k}\{\textbf{C}\rho_{k}+(1-\rho_{k})\textbf{I}^{m}\}$ and $\rho_{k}=\gamma_{k}^{2}/(\gamma^{2}_{k}+\delta_{k}^{2})$ is the phylogenetic signal parameter per covariate (equivalent to Pagel’s $\lambda$, Pagel 1999; Pearse, Davies, and Wolkovich 2023), while $\sigma^{2}_{k}=\gamma^{2}_{k}+\delta^{2}_{k}$ is the variance of the sum of the two terms. Consequently, the variances of the original two terms can be retrieved from this alternative model formulation by nothing that $\gamma^{2}_{k}=\rho_{k}\sigma_{k}^{2}$ and $\delta^{2}_{k}=(1-\rho_{k})\sigma^{2}_{k}$. Equivalently, $\boldsymbol{\beta}_{k}^{\epsilon}\sim\mathcal{N}(\textbf{0},\boldsymbol{\Sigma}_{k})$. The latter formulation has the benefit of only including one set of random effects that needs to be integrated out of the likelihood (which is usually computationally intensive for many random effects), but the downside that it requires explicit inversion of the covariance matrix for species at each iteration during model fitting (which tends to be computationally intensive for many species).
Covariances between the random effects of different covariates $\boldsymbol{\Sigma}^{r}$, i.e., for the rows of $\textbf{B}^{\epsilon}$ can be introduced by instead defining the distribution of the random effects over all species and covariates simultaneously. Let $\textbf{L}(\boldsymbol{\Sigma}_{k})$ denote the lower cholesky factor of the covariance matrix for the $k^{\text{th}}$ covariate effect, so that instead we parameterize the covariance over the random effects for all species and covariates as
$\text{vec}(\textbf{B}^{\epsilon\top})\sim\mathcal{N}\{\textbf{0},\textbf{L}(\boldsymbol{\Sigma}^{r}\otimes\textbf{I}^{m})\textbf{L}^{\top}\},$ (2)
where L is a block-diagonal matrix with $\textbf{L}(\boldsymbol{\Sigma}_{1}),\dots,\textbf{L}(\boldsymbol{\Sigma}_{p})$ as block matrices. In this formulation, $\boldsymbol{\Sigma}^{r}$ controls correlation between covariate effects of the same species, and together with the phylogenetic signal parameters also the correlation between covariate effects of different species. These off-diagonal blocks in the covariance matrix are by extension formulated as a lower-upper (LU) decomposition and take the form $\textbf{L}(\boldsymbol{\Sigma}_{k})\textbf{L}(\boldsymbol{\Sigma}_{l})^{\top}\Sigma_{kl}$, for covariates $k,l=1\ldots p$. The lower triangular matrix of this LU decomposition is $\textbf{L}(\boldsymbol{\Sigma}_{k})\text{Diag}[\text{diag}\{\textbf{L}(\boldsymbol{\Sigma}_{k})\}^{-1}$, where $\text{Diag}(\cdot)$ constructs a diagonal matrix from a vector and $\text{diag}(\cdot)$ extracts the diagonal of a matrix. The associated upper triangular matrix is then $\text{Diag}[\text{diag}\{\textbf{L}(\boldsymbol{\Sigma}_{k})\}]\textbf{L}(\boldsymbol{\Sigma}_{l})^{\top}\Sigma_{kl}$. If instead we assume $\rho_{k}=\rho$, i.e., that the phylogenetic signal is the same for all covariates, equation (2) reduces to equation (4) of Ovaskainen et al. (2017) or equation (11.1-2) of de Villemereuil and Nakagawa (2014), i.e., $\text{vec}(\textbf{B}^{\epsilon})\sim\mathcal{N}(\textbf{0},\boldsymbol{\Sigma}\otimes\boldsymbol{\Sigma}^{r})$.
Consequently, the matrix of species associations at sites $i$ and $i2$ due to the phylogenetic mixed-effects model with multiple signal parameters is:
$\boldsymbol{\Sigma}_{spp,i,i2}=(\textbf{x}_{i}^{\top}\otimes\textbf{I}^{m})\textbf{L}(\boldsymbol{\Sigma}^{r}\otimes\textbf{I}^{m})\textbf{L}^{\top}(\textbf{x}_{i2}\otimes\textbf{I}^{m}).$
The model in equation (1) can be extended in various ways, for example by incorporating unconstrained latent variables as in Niku et al. (2021), or constrained or informed latent variables (van der Veen et al. 2023), potentially with a quadratic response model (van der Veen et al. 2021), with random site effects to account for pseudoreplication (as included in the following case study) or spatial autocorrelation, or with additional fixed effects, but we have omitted those here for brevity.
Similarly, in its present form the model only incorporates positive species associations. It is technically possible to adjust the model to additionally incorporate negative species assocations, but we consider such an implementation as a potential future avenue for research (see appendix S1 for more information).
Parameter estimation
In this section we elaborate on the four approximations that are implemented to quickly fit the phylogenetic random effect models. We evaluate the accuracy of these stacked approximations using simulation studies in the following section.
For brevity, formulating the model in terms of a single set of random effects with covariance matrix $\boldsymbol{\Sigma}$ given in equation (2). With $f(y_{ij}|\boldsymbol{\beta}^{\epsilon}_{j})$ a response distribution such as a Tweedie or negative binomial distribution, and $\Theta$ a vector incorporating of all other parameters in the model (such as scale parameters), we have the following marginal log-likelihood function:
$\mathcal{L}(\boldsymbol{\Theta})=\sum\limits_{i=1}^{n}\log\biggl{[}\int\prod\limits_{j=1}^{m}f\biggl{\{}y_{ij}|\boldsymbol{\beta}^{\epsilon}_{j}\biggr{\}}\mathcal{N}\biggl{\{}\text{vec}\biggl{(}\textbf{B}^{\epsilon}\biggr{)};\boldsymbol{\Sigma}\biggr{\}}d\textbf{B}^{\epsilon}\biggr{]}.$ (3)
Variational Approximations
To efficiently retrieve a closed-form approximation to the log-likelihood in equation (3), we use Variational Approximations (VA, Hui et al. 2017). VA introduces hyperparameters (“variational parameters”) that need to be optimised in order to improve the approximation (tightening the lower bound). With many species and random effects this can significantly slow down model fitting. Other methods for fitting generalized mixed-effects models, such as the Laplace approximation, suffer from similar issues with high dimensions. The Laplace approximation requires evaluating the matrix of second derivatives of the joint likelihood for the random effects and the data, which also significantly slows down for a large number of (correlated) random effects. VA has the benefit of allowing for a second approximation by introducing a more simple structure for the VA covariance matrix, so that the size of the optimisation problem is reduced significantly for quicker model fitting.
Various VA covariance structures are possible to formulate, but some require stronger assumptions (such as independence of species or random effects), or have a considerably higher number of variational parameters (e.g., as in an unstructured covariance matrix), so that they require significantly longer computational times. In one of the most reduced scenarios, we assume the variational distribution for the random effects to follow a matrix normal distribution, i.e., $q(\textbf{B}^{\epsilon})=\mathcal{MN}(\textbf{a},\textbf{A}^{r},\textbf{A}^{m})$ with a a $p\times m$ matrix of variational mean parameters, $\textbf{A}^{r}$ a $p\times p$ VA covariance matrix, and $\textbf{A}^{m}$ a $m\times m$ VA covariance matrix. Some different formulations are provided in the supplementary information (appendix S2) and implemented in the gllvm R-package. For the remainder of this article, we limit ourselves to application of the matrix normal case, with unstructured or diagonal $\textbf{A}^{r}$. We show in the simulations below that the choice of a matrix normal VA distribution still enjoys high accuracy, despite its relative inflexibility, even with diagonal $\textbf{A}^{r}$.
With this choice of a matrix normal covariance structure, VA can still be slow due to the potentially large number of variational parameters in $\textbf{A}^{m}$. The number of variational parameters in this matrix grows quadratically in the number of species. To ensure better scaling, and quick computation also with many species, we employ a third approximation by representing the off-diagonal entries of $\textbf{A}^{m}$ in a reduced-rank fashion so that $d<<m$, i.e., $\textbf{A}^{m}=\textbf{A}^{d}\textbf{A}^{d\top}+\textbf{A}^{m-d}$ where $\textbf{A}^{d}$ is a $m\times d$ lower triangular matrix with positive diagonal entries, and $\textbf{A}^{m-d}$ is a $m\times m$ diagonal matrix with the first $d$ diagonal entries equal to zero and the remaining entries positive, ensuring that $\textbf{A}^{m}$ is of full rank. Such a choice for reducing the number of variational parameters in the likelihood is optimal for computation speed, to reduce memory use during model compilation, and to reduce the size of the Hessian matrix that needs to be inverted for parameter estimation and for post-hoc estimating (approximate) statistical uncertainties. This does introduce the additional problem of having to select the rank of the variational covariance matrix, but in the simulation studies that follow we show that high accuracy can be achieved even when using only a single dimension.
Sparse precision matrix
The final remaining computational bottleneck involves inversion of the covariance matrix for the random effects. Since the covariance matrix of the covariate effects $\boldsymbol{\Sigma}^{r}$ is usually relatively small, its inverse can be evaluated explicitly without excessive computational cost, and its determinant straightforwardly retrieved by formulating it following a log-cholesky parameterization (Pinheiro and Bates 1996). In contrast, the number of species in multivariate ecological data is often relatively large, so that $\boldsymbol{\Sigma}^{-1}_{k}$ is expensive to calculate. With a single phylogenetic signal parameter over all covariate effects, the form of the covariance function for $\boldsymbol{\Sigma}_{k}$ allows inversion by repeatedly applying the Sherman-Morrison formula for computationally cheap rank-one updates of the form $\textbf{C}^{-1}=\textbf{C}^{-1}-\{1-\rho\}/\{1+(1-\rho)\textbf{C}_{j}\textbf{C}_{j}^{\top}\}$, but further improvement is warranted.
Let $A_{j}$ a set of indices (species) used to condition a random effect on. Consequently, the joint distribution for the random effect over all species can be written as $p(\boldsymbol{\beta}^{\epsilon}_{k})=\prod\limits^{j=1}p(\beta^{\epsilon}_{jk}|\{\beta^{\epsilon}_{1k},\beta^{\epsilon}_{2k},\ldots,\beta^{\epsilon}_{j-1k}\}:A_{j})$, and while momentarily ignoring potential correlations between random effects, each conditional distribution can be written as $p(\beta^{\epsilon}_{jk}|\{\beta^{\epsilon}_{1k},\beta^{\epsilon}_{2l},\ldots,\beta^{\epsilon}_{j-1k}\}:A_{j})=\mathcal{N}(\boldsymbol{\Sigma}_{kjA_{j}}\boldsymbol{\Sigma}_{kA_{j}A_{j}}^{-1}\boldsymbol{\beta}^{\epsilon}_{A_{j}k},\boldsymbol{\Sigma}_{kjj}-\boldsymbol{\Sigma}_{kjA_{j}}\boldsymbol{\Sigma}_{kA_{j}A_{j}}^{-1}\boldsymbol{\Sigma}_{kA_{j}j})$.
Wist and Rue (2006) describe various methods for specifying Gaussian Markov Random Fields with a sparse precision matrix, such as by thresholding or by a predefined sparsity pattern (such as a band pattern). Here, we mostly focus on adopting NNGPs as our fourth and final approximation, but compare with approximating the precision matrix by a band pattern in the real data example. Unlike regular matrix inversion, that tends to be inherently sequential (such as repeated application of the Sherman-Morrison formula), NNGPs facilitate parallel computation well because the calculation can be performed for all random effects separately. In NNGPs we construct the (approximate) inverse of the cholesky factor of the covariance matrix, which makes for an efficient implementation of equation (2), and combines nicely with the likelihood in equation (3).
NNGPs are usually applied in a spatial (and Bayesian) context (Finley et al. 2019), truncating the set of indices $A_{j}$ for spatial locations based on geographical proximity. For a phylogenetic random effect we instead condition each species’ random effect on the species closest in evolutionary space. Distance between species in evolutionary space is readily accessible from the phylogeny, and calculated as the total shared branch length on the phylogeny. It is this truncation for the conditional distributions of the random effects that implies a sparse precision matrix for the joint distribution of the random effects. Note, NNGPs are order-sensitive, so that the approximation can suffer from a poor ordering of species. However, similarly to Guinness (2018), we choose to see this as a feature of the methodology, so that the approximation can be improved by permuting the columns of the response data. We further explore this issue of order sensitivity in the real data example.
Simulation studies
To evaluate the accuracy of the proposed combination of approximations, we performed various simulation studies, but for brevity we assumed $\rho_{k}=\rho$, i.e., a single phylogenetic signal parameter for all covariates. We performed simulations separately to evaluate the accuracy of the low-rank VA approximation, and of the NNGP approximation. To do so, we simulated datasets following equation (1), without functional traits, and while assuming that the phylogenetic signal parameter is the same for all covariates. We focus on estimation accuracy of the phylogenetic signal, as it conveniently provides a single measure to assess, but also because we deem it the most likely quantity in the model to be affected by changes in the approximation.
For the first simulation study, we kept the number of sites fixed at $n=100$ and the number of species at $m=200$. We simulated 250 datasets while varying the rank of the VA covariance matrix ($d\in\{1,3,5,10,25,100,200\}$). Instead of the NNGP, in this first simulation study we directly calculated the inverse of the phylogenetic covariance, in order to avoid introducing potential error due to the NNGP and so that we may isolate the error solely due to the matrix normal VA approximation with reduced rank, and unstructured or diagonal $\textbf{A}^{r}$. For comparison we additionally fitted the model with the Laplace approximation as automatically applied by Template Model Builder (TMB, Kristensen et al. 2015). We re-simulated the phylogeny for each simulation using the ape R-package (Paradis and Schliep 2019). We simulated five covariates from $\text{Uniform}(-1,1)$, the covariance matrix for the random effects from $W(\frac{1}{5}\textbf{I}^{p},p)$, and kept $\rho=0.5$ fixed so that the median absolute error is at best zero and at most 0.5. As the total number of simulations is high, and the models take long to fit when the rank of the VA approximation is large, we only simulated binary datasets. Binary datasets contain least information of the datatypes that are accommodated by the gllvm R-package, so that these simulations should provide a reliable upper bound to the accuracy of the approximation. The results of the simulations are summarized in Figure 1.
Figure 1: Results of the first simulation study for evaluating the accuracy of the low-rank VA approximation. The left panel displays increase of fitting time with the rank (1, 3, 5, 10, 25, 100, 200) of the VA approximation, with solid black lines indicating models fitted with the VA approximation and with unstructured $\textbf{A}^{r}$ (circles) or diagonal $\textbf{A}^{r}$ (triangles), while the dashed red line indicates the computation time for the models fitted by Laplace approximation. The faded black lines represent the 2.5% and 97.5% percentiles for the VA approximation and the faded dashed red lines for the Laplace approximation. The right panel shows the median absolute error for the phylogenetic signal parameter, which at most can be 0.5.
The first simulation study confirms that the computation time of the models rapidly increase with the selected rank of $\textbf{A}^{m}$ in the VA approximation, while being competitive to the Laplace approximation implemented with TMB (indicated by the red dashed line). As seen in the left panel of Figure 1, the rank of the VA approximation has a large impact of the computation time, but fortunately little impact on the estimation accuracy of the phylogenetic signal parameter, which remained nearly stable across various ranks of the VA approximation and was indeed lowest for the full rank approximation. Models fitted with an unstructured $\textbf{A}^{r}$ in the VA approximation exhibited longer fitting times, so that it took longer to fit models than with the Laplace approximation. Models fitted with a diagonal $\textbf{A}^{r}$ fitted more quickly than with the Laplace approximation, while also providing more accurate estimates for the phylogenetic signal than with an unstructured $\textbf{A}^{r}$.
In the second simulation study we kept the rank of the VA approximation fixed to one and $\textbf{A}^{r}$ unstructured, and fixed the number of sites to $n=100$, but instead varied the number of species ($m\in\{50,100,250,500\}$) and the number of nearest neighbors used ($nn\in\{1,3,5,7,9,11,13,15\}$) to approximate the inverse of the phylogenetic covariance matrix. For each combination we simulated 250 datasets, for either binary data or counts. The results are presented in Figure 2.
Figure 2: Results for the second simulation study to test the ability of the models to accurately estimate the phylogenetic signal parameter using the low-rank VA and NNGP approximations for different numbers of species $m$ in the data.
The results in Figure 2 represent the compounded approximation error due to both the low-rank VA approximation and the NNGP. The number of species has a larger impact on the estimation accuracy of the phylogenetic signal than the NNGP approximation. As the number of species increases, the estimation accuracy of the phylogenetic signal improves, with $m=50$ potentially requiring a much larger number of nearest neighbouurs for accurate estimation, but with $m>100$ providing accurate estimates under most scenarios. As expected, the accuracy was higher with more nearest neighbours, but a relatively low (e.g., 10) number of nearest neighbors seems sufficient for reasonably accurate estimates of the phylogenetic signal, even for large numbers of species in the data. Though not presented here, application of NNGPs reduced computation time significantly relative to the models fitted for the first simulation study.
Case study
We demonstrate the new method for fitting phylogenetic random effect models on data collected by Abrego et al. (2022a). The dataset contains the binary responses of 320 wood-inhabiting fungi surveyed on 1809 logs across 53 European beech (Fagus sylvatica) forest sites in different European countries that were grouped into eight regions, so that there were in total $n=1666$ rows in the data. Species with four or less presences were removed before analysis in the original article. Abrego et al. (2022a) fitted JSDMs with the Hmsc R-package with as main goal to determine if traits and phylogenetic relationships structure species communities differently at different spatial scales. To that end, their analyses included environmental variables both at the log- and site-scales such as: diameter at breast height of logs (DBH), decay stage of the logs (on a scale of one to five), an index for connectivity (at the 10km scale), area of forest patches (in hectares), annual temperature range (degrees Celsius) and annual precipitation, and naturally a phylogeny.
The models fitted by Abrego et al. (2022a) additionally incorporated six latent variables at the region-level, and although the gllvm R-package facilitates including such latent variables at the group-level, we have here chosen to omit the latent variables for brevity. In the Hmsc R-package each model took about a week to 10 days to run (N. Abrego, pers. comm., 1st March 2024). Results for the same four models as in are included in appendix S3. It takes about six to twenty minutes to fit a model with covariate-specific phylogenetic signal parameters (on a Dell Latitude 7490 using 7 CPU, with $nn=15$ neighbours, and with unstructured $\textbf{A}^{r}$), and two to four minutes for models while assuming $\rho_{k}=\rho$, i.e., a single phylogenetic parameter for all covariate effects. Fitting the models with diagonal $\textbf{A}^{r}$ can further reduce computation time, to (e.g.) 1.5 minute (corresponding to the two minutes for a model without traits). These models incorporated nested random intercepts for the study region and reserves, and including all correlation parameters between the covariate effects.
As previously noted, the quality of the NNGP approximation tends to be sensitive to the order of species. This has previously been noted by Guinness (2018) in a spatial context, who also explored orderings for improving the NNGP approximation. Similarly to finding permutation methods for reducing fill-in of the cholesky factor of a sparse matrix (Wist and Rue 2006), identifying an optimal ordering for NNGPs is an NP-complete problem. Consequently, heuristic methods are essential for determining an ordering that improves the NNGP approximation.
The approximation error for the phylogenetic precision matrix will have to be assessed on a case-by-case basis, but can be explored before model fitting, because the approximation error for $\boldsymbol{\Sigma}_{k}$ is clearly related to that of C. However, the former is not available before model fitting, whereas the approximation to $\textbf{C}^{-1}$ is. We define the approximation error as $\|\widetilde{\textbf{C}^{-1}}\textbf{C}-\textbf{I}^{m}\|_{F}$, where $\widetilde{\textbf{C}^{-1}}$ denotes the approximation to $\textbf{C}^{-1}$.
Rather than extensively searching for potentially suitable orderings, we instead focus on exploring the approximation error due to six orderings that are easily accessible to ecologists: 1) the order of species in the phylogeny; because species with large covariances are naturally located closely together in the phylogeny, 2) alphabetically; as a baseline and because it is how data is often provided in practice, 3) the sum of pairwise distances of species, 4) the distance of species to the root of the phylogeny, 5) by the first eigenvector of C (as in Matsuba et al. 2024), 6) by the sum of squared covariances. We additionally explore the approximation error using a band pattern (as in Wist and Rue 2006), ignoring the distance between species in the phylogeny. The results are visualized in Figure 3.
Figure 3: Approximation error for the band pattern approximation and the NNGP, for different permutations of species in the fungi data.
The results show that the NNGP has the potential of a smaller approximation error, though it requires conditioning on six or more species. When conditioning on less than six species, the pre-defined band pattern approximation exhibits a smaller approximation error. For the band pattern approximation, ordering the response data following the order of species in the phylogeny provides the smallest approximation error, while for the NNGP this is when ordering the response data by the total shared branch length in the phylogeny. In both instances, ordering species alphabetically made for the worst approximation, even though as a result species within the same genus will usually be located in close proximity (which can be expected to be phylogenetically similar, and thus have large covariances). The same does not hold for genera that are phylogenetically similar.
Figure 4: The phylogenetic covariance matrix as a result from the sparse approximation to the precision matrix for different orderings and numbers of nearest neighbours (nn). Three orderings are displayed, corresponding to poor (alphabetical), average (tips) and good (distance) approximations.
The approximation error affects both the estimation accuracy and computation speed. To demonstrate this, we fitted three models to the example data using three different orderings and $nn\in\{5,10,15,20\}$, the results are displayed in Table 1. The models included three trait covariates, all environmental covariates, random intercepts for region and reserve. We also assumed a diagonal $\textbf{A}^{r}$ for faster fitting. In Figure 5 the community mean responses and species’ random effects for the model with $nn=20$ with distance ordering are displayed alongside the phylogeny.
Table 1: Phylogenetic signal parameter estimates for models with different orderings of species in the response data, and number of nearest neighbours (nn), fitted to the Abrego et al. (2022) data. A parameter estimate of 0.05 or lower is displayed as a blank cell for improved readability.
Order nn Time (min) Intercept Deadwood size Decay stage Decay stage² Connectivity Temperature Precipitation ln(Reserve area)
alphabetical 5 8.77 0.13 1.00 0.88 0.13 0.51 0.55
alphabetical 10 5.92 0.24 0.65 1.00 0.87 0.10 0.74 0.68
alphabetical 15 5.96 0.22 0.66 1.00 0.88 0.14 0.71 0.70
alphabetical 20 9.33 0.22 0.64 1.00 0.88 0.15 0.64 0.71
tips 5 11.98 0.17 0.62 0.99 0.89 0.10
tips 10 6.96 0.07 0.98 0.89 0.52 0.74
tips 15 7.81 0.49 0.99 0.90 0.59 0.77
tips 20 8.07 0.50 0.99 0.91 0.59 0.81 0.11
distance 5 5.63 0.18 0.68 0.88 0.12 0.10 0.16
distance 10 6.02 0.26 0.99 0.86 0.07
distance 15 8.74 0.31 0.99 0.85
distance 20 6.07 0.31 0.98 0.87 0.19 0.46
Figure 5: The right panel displays community mean responses to the covariates, with 95% confidence intervals. Coefficients are greyed out if the confidence interval crosses zero, so that it is unclear if there is an effect of a covariate on the community. The bottom panel presents species-specific deviations from the community mean effects. A species-specific effect of zero is interpreted as lack of evidence that a species responds differently to the environment than the community mean response. Orange colors indicate a more negative response than the community mean effect, and green more positive. Species effects of which the prediction interval included zero have been removed.
Discussion
In this article, we developed a new method for quickly fitting phylogenetic mixed-effects models, akin to the approaches in the phyr (Ives et al. 2020), Hmsc (Tikhonov et al. 2024), and MCMCglmm (Hadfield 2010) R-packages. Because our models are developed as part of the gllvm R-package (Niku et al. 2023), the software supports a wide range of response types, including for Bernoulli, (zero-inflated) Poisson and negative-binomial, Tweedie, and beta responses (Korhonen et al. 2021; Niku et al. 2017; Hui et al. 2017).
By utilizing approximate likelihood methods, the models are magnitudes faster to fit than in the Hmsc and MCMCglmm R-packages, which are based on Markov chain Monte Carlo methods. By stacking approximations, namely a VA approximation with matrix normal covariance structure, combined with low-rank approximation of one of the matrices, and NNGPs (Finley et al. 2019) for calculating a sparse approximation to the phylogenetic precision matrix, computation time of the proposed models is reduced as much as possible. Two sets of simulation studies confirm that the proposed combination of approximations still enjoys accurate estimation of the phylogenetic signal, while significantly speeding up model fitting.
We purposefully limited the flexibility of the VA approximation to fit phylogenetic random effects models more quickly. The flexibility of VA is determined by the number of hyperparameters introduced to the likelihood, which are estimated jointly with the model parameters. Especially for VA, we find that a good optimisation routine, besides decent initial values, is critical for successful and speedy convergence of the models. In our experience, limited-memory optimisation algorithms (Nocedal and Wright 2006) often lead to faster model fitting and convergence, and an improved ability to successfully leverage methods for parallel computation. In future research we hope to explore improved optimisation routines (as in van der Veen 2024; Kanzow and Steck 2023), and their ability to more efficiently fit mixed-effects models. In our simulation study the VA approximation was competitive in computational speed to the Laplace approximation as applied using Template Model Builder (Kristensen et al. 2015), though this was limited to 200 species and five covariates, while we expect the VA approximation to especially outperform the Laplace approximation with a higher number of covariates and species. When reducing the number of parameters in the VA approximation as much as possible, the method fits faster than when models are fitted using the Laplace approximation.
The cost of quickly fitting the models comes in form of selecting the number of nearest neighbours for the NNGP approximation. The models are sensitive to the number of neighbours, which thus needs to be carefully assess prior to model fitting. Fortunately, this can straightforwardly be done by quantifying the approximation error under different orderings of species in the data, and numbers of nearest neighbours, and by picking the best combination. We demonstrated picking the number of nearest neighbours and species’ ordering using real data on 215 species of wood-inhabiting fungi (Abrego et al. 2022b). One of the main novelties of the proposed approach for fitting phylogenetic random effects models is the ability to estimate phylogenetic signal separately for each environmental covariate, without the need to introduce an additional set of random effects as in the phyr R-package, and while simultaneously facilitating the estimation of correlation between random effects for different environmental covariates. Abrego et al. (2022a) fitted various models with and without trait covariates to determine if phylogenetic structuring of species responses occurred at multiple spatial scales simultaneously: at the forest site-scale and at the scale of deadwood logs. Abrego et al. (2022a) post-hoc estimated covariate-specific phylogenetic signal, while here we estimated the covariate-specific phylogenetic signal directly as part of the model, and while fitting the models in minutes, rather than days as in the Hmsc R-package.
The phylogenetic signal parameter represents the degree to which species environmental responses are phylogenetically structured, and the associated variance of covariates the rate at which the responses evolve. When the phylogenetic signal parameter is one, species responses to an environmental covariate are fully phylogenetically structured, and at zero the responses are independent (Pearse, Davies, and Wolkovich 2023). The reason for the presence of phylogenetic signal, or the lack thereof, cannot be discerned from phylogenetic random effects models directly. When there is phylogenetic signal, this can be due to the lack of (phylogenetically structured) continuous trait covariates in the model, because the measured trait covariates are not phylogentically structured, or because they are structured in a way that disagrees with the structuring in species environmental responses. In contrast, a lack of phylogenetic signal may be because the trait covariates in the model sufficiently introduce phylogenetic structuring to the species responses, because there is no ongoing evolution in the trait that is species environmental responses, or because evolution moves so rapidly that closely related species are (apparently) so different that their environmental responses are independent.
Various avenues for future research exist to extend the model. Here, we assumed the covariance matrix for the covariate effects to be unstructured. There are various scenarios where a structured covariance matrix would me more sensible, such as for phylogenetically structured site intercepts. Sites are not usually randomly distributed in space or time, while structuring the random effects with an autoregressive or spatial autocorrelation function (e.g., as in Brooks et al. 2017) could address such issues. For computational reasons, such an extension would require an additional approximation to the spatial covariance matrix (similar to the methods used in this article). Although technically possible, combining (control of) further approximations with an intuitive interface might prove challenging.
In their current form phylogenetic random effect models only incorporate positive species associations, meaning that more closely related species can only occur in more similar environments. This makes sense if closely related species segregate their resource utilization. However, two closely related species that share the same resource would be more likely to occur in different environments, i.e., exhibit a negative association. Phylogenetic repulsion can be modelled in the current framework, by prior to model fitting replacing the phylogenetic correlation matrix by its inverse (Ives and Helmus 2011). Still, the current model is limited to either attraction or repulsion, and ideally its form should be adjusted to allow the data to drive that decision instead.
Acknowledgements
The authors thank Sara Taskinen for comments on an earlier version of the manuscript.
Data availability
The data used in this study is available online (Abrego et al. 2022b).